# A Framework for Double-Blind Federated Adaptation of Foundation Models

### v1

@@ 4.3 Local Learning
Brute-force as the optimal attack strategy: Given the non-uniqueness property above, an attacker aiming to recover ${\bm{A}}$, ${\bm{B}}$, and ${\bm{C}}$ must resort to brute force, testing all possible configurations of $({\bm{A}},{\bm{B}},{\bm{C}})$ that satisfy the products $({\bm{P}},{\bm{Q}},{\bm{R}})$. The total number of combinations of $({\bm{A}},{\bm{B}},{\bm{C}})$ the attacker would need to test is $n!$ and for $n=16$, we need $16!\approx 2\cdot 10^{13}$ combinations.
4.4 Secure Aggregation
To ensure the secure aggregation of model parameter updates from clients, we leverage secure multi-party computation (MPC) for computing averages in a privacy-preserving manner. This approach, commonly referred to in the literature as Secure Aggregation [3], enables the aggregation server to compute the average of client updates without gaining access to the individual updates themselves.
Datasets Methods Centralized Federated $(K=5)$
CIFAR-10 Linear probing $0.9226$ $0.9203$ $0.9191$ $0.7447$
Full fine-tuning $0.9635$ $0.9759$ $0.9725$ $0.8857$
Ours $0.9428$ $0.9471$ $0.9413$ $0.8540$
CIFAR-100 Linear probing $0.7476$ $0.7486$ $0.7414$ $0.5317$
Full fine-tuning $0.8361$ $0.8684$ $0.8611$ $0.7882$
Ours $0.7930$ $0.7929$ $0.7808$ $0.6620$
SVHN Linear probing $0.5938$ $0.5879$ $0.5732$ $0.3385$
Full fine-tuning $0.9680$ $0.9763$ $0.9692$ $0.7601$
Ours $0.9232$ $0.9329$ $0.9249$ $0.7431$
Table 1: Comparison of the performance of our method against baseline approaches across three datasets (CIFAR-10, CIFAR-100, and SVHN) in both centralized and federated learning settings. The federated experiments involve five clients $(K=5)$ with data partitioned using a Dirichlet distribution at varying levels of heterogeneity $(\alpha=100,1,0.01)$.
5 Experiments and Results
5.1 Datasets
To validate and implement our approach, we will utilize a well-known vision transformer (ViT) pretrained on ImageNet-1K (ViT-Base) as the FM. The public auxiliary datasets for distilling the FHE-friendly FM are:
  • Tiny-Imagenet [26] is a subset of the larger ImageNet dataset, and it contains 200 different classes, each with 500 images, 100 validation/test images, totaling 120000 images.
  • Fed-ISIC2019 [50] is a multi-class dataset of dermoscopy images containing $23247$ images across $8$ melanoma classes, with high label imbalance. This dataset provides a valuable opportunity to test our framework within the healthcare domain. For the distillation phase of our experiments, we exclusively use data from client 1.
The private downstream datasets are:
  • CIFAR-10 and CIFAR-100 [25] datasets are standard benchmarks for image classification, containing $60000$ color images across $10$ and $100$ classes, respectively.
  • SVHN [38] dataset is a benchmark for digit recognition, consisting of over $600000$ color images of house numbers. We utilize $73257$ images for training and $26032$ for testing, disregarding the remainder.
  • Fed-ISIC2019. The remaining data points for centers 2-6 are used in the fine-tuning experiments, aligning well with the federated setup, as the dataset is tailored for federated learning. For the centralized setup, all data points are pooled into a single client.
5.2 Experimental Setup
Public dataset Methods Centralized Federated
Fed-ISIC2019
(center=0)
Linear probing $0.6599$ $0.5856$
Full fine-tuning $0.7811$ $0.6752$
Ours $0.7090$ $0.6679$
Tiny-Imagenet
Linear probing $0.6372$ $0.5789$
Full fine-tuning $0.7817$ $0.6985$
Ours $0.7051$ $0.6481$
Table 2: Performance comparison of our method with baseline approaches on the Fed-ISIC2019 dataset with five clients ($K=5$), using two auxiliary datasets: Fed-ISIC2019 (center=0) as an in-distribution (InD) dataset and Tiny-ImageNet as an out-of-distribution (OOD) dataset.
5.3 Results
Figure 4: Accuracy vs. no. of trainable parameters trade-off (x-axis in log scale) derived from Table 1, illustrating the performance of the methods across three datasets.
Main results: In Table 1, the performance of our proposed method and baseline methods across three datasets (CIFAR-10, CIFAR-100, and SVHN) is evaluated under both centralized and federated learning settings. The federated learning experiment uses a Dirichlet partitioning strategy with varying levels of data heterogeneity, controlled by the Dirichlet concentration parameter $\alpha$ (ranging from $100$ to $0.01$). The results demonstrate that full fine-tuning achieves the highest accuracy across all datasets and settings, particularly excelling in more homogeneous federated scenarios, but it is computationally expensive. Linear probing performs well under centralized and homogeneous federated settings. When the task is relatively more straightforward, it fails to work with the SVHN dataset, reaching only $0.5938$ accuracy in the centralized setting, and struggles when there is extreme heterogeneity. Our proposed method delivers robust and competitive results across all settings, showing strong resilience to non-i.i.d data, making it suitable for practical federated learning scenarios (refer to Figure 4).
Fed-ISIC2019 results: Table 2 compares the performance of our method and baselines on the Fed-ISIC2019 dataset with five centers. The auxiliary datasets used at the LSP include (1) Fed-ISIC2019 with only the first center (treated as an in-distribution dataset) and (2) Tiny-ImageNet (treated as an out-of-distribution dataset). The results demonstrate that knowledge transfer from the OOD dataset is effective for all the methods, highlighting that the public dataset can be any available dataset. Notably, our method effectively learns from the OOD FM, with a minimal drop compared to the InD FM, highlighting its robustness.
Methods $K=10$ $K=20$ $K=50$
Linear probing $0.9167$ $0.9142$ $0.9007$
Full fine-tuning $0.9739$ $0.9513$ $N/A$
Ours $0.9446$ $0.9422$ $0.9287$
Table 3: Scalability analysis of the proposed method to baseline approaches on the CIFAR-10, with varying number of clients $K\in\{10,20,50\}$ under a Dirichlet concentration parameter of $1.0$ for data partitioning. $N/A$- one GPU is insufficient to run the experiment.
Scalability results: Table 3 illustrates the scalability of our method and the baseline approaches on the CIFAR-10 dataset with increasing number of clients $(N=\{10,20,50\})$ using Dirichlet partitioning $(\alpha=1.0)$ and a fixed batch size of 8. Full fine-tuning achieves the highest accuracy for $K=10$ and $K=20$ but becomes infeasible for $K=50$ due to GPU limitations (we use only one GPU for each of our experiments). Linear probing demonstrates stable performance, but our method outperforms linear probing in all the settings, balancing compute efficiency, scalability, and accuracy and demonstrating its suitability for federated setups with a large number of clients.
6 Conclusion
This paper has explored the synergistic integration of FL and FM, which offers a promising framework for advancing AI technologies while adhering to data privacy and security standards. FL provides significant advantages in data privacy, distributed model training, and scalability. Concurrently, FM contributes with its pre-trained capabilities and robust performance across various domains. Our study has particularly concentrated on the efficient fine-tuning of large FMs within a federated framework. We addressed the critical challenges associated with the memory, runtime, and computational demands of deploying large FMs, along with ensuring data privacy, confidentiality, and system security in decentralized environments.
Supplementary Material
Appendix A How Does the Distillation Work?
@@
We adopt the transformer distillation approach [21], that consists of two key phases: (i) transformer-layer distillation and (ii) prediction-layer distillation. For generality, let the original transformer model be referred to as the teacher model $(\mathcal{T})$ and the approximation-enabled transformer model as the student model $(\mathcal{S})$. We assume that both models share the same architecture but differ only in their non-linear components (Softmax, GELU, LayerNorm). Further, we outline the primary sub-layers of the transformer blocks:
@@
where $h$ is the number of attention heads. And, the distillation of the hidden representation is formulated as follows:
@@
=$\mathcal{L}=\begin{cases}\mathcal{L}_{a}+\mathcal{L}_{h}&\qquad\triangleright\ \text{Stage I},\\ \mathcal{L}_{p}&\qquad\triangleright\ \text{Stage II}.\end{cases}$ (16)
@@
Figure 5: Schematic illustration of the layer-wise knowledge distillation (Equations 12 and 13).
Appendix B Experimental Results
Datasets Methods No. of params (M) Latency (ms) Memory (GB)
CIFAR-10 / SVHN Linear probing $0.0073$ (1.00×)(1.00\ $\times$)( 1.00 × ) $15.1$ (1.00×)(1.00\ $\times$)( 1.00 × ) $9.39$
Full fine-tuning $82.1096$ (11247.89×)(11247.89\ $\times$)( 11247.89 × ) $47.4$ (3.14×)(3.14\ $\times$)( 3.14 × ) $18.13$
Ours $0.2536$ (34.74×)(34.74\ $\times$)( 34.74 × ) $15.7$ (1.04×)(1.04\ $\times$)( 1.04 × ) $9.08$
CIFAR-100 Linear probing $0.0733$ (1.00×)(1.00\ $\times$)( 1.00 × ) $15.7$ (1.00×)(1.00\ $\times$)( 1.00 × ) $9.40$
Full fine-tuning $82.1756$ (1121.09×)(1121.09\ $\times$)( 1121.09 × ) $47.4$ (3.03×)(3.03\ $\times$)( 3.03 × ) $18.13$
Ours $0.3196$ (4.36×)(4.36\ $\times$)( 4.36 × ) $16.3$ (1.04×)(1.04\ $\times$)( 1.04 × ) $9.08$
Fed-ISIC2019 Linear probing $0.0059$ (1.00×)(1.00\ $\times$)( 1.00 × ) $20.2$ (1.00×)(1.00\ $\times$)( 1.00 × ) $17.32$
Full fine-tuning $82.1082$ (13916.64×)(13916.64\ $\times$)( 13916.64 × ) $51.2$ (2.54×)(2.54\ $\times$)( 2.54 × ) $8.71$
Ours $0.2522$ (42.75×)(42.75\ $\times$)( 42.75 × ) $21.1$ (1.05×)(1.05\ $\times$)( 1.05 × ) $8.39$
B.1 Main Results
Table 4 provides additional details on the number of trainable parameters, the latency, and the GPU memory required to execute the respective experiments across the given datasets and methods. These results complement Tables 1 and 2. As shown in the table, our proposed method closely matches Linear Probing in terms of the number of parameters and the latency required to process a single sample, while full fine-tuning exhibits a significant disparity in effectiveness. Additionally, the GPU memory required for our method is comparable to that of Linear Probing, whereas full fine-tuning demands nearly twice as much memory.
B.2 Distillation Results
Following the findings discussed in Section A, we present performance plots for two key processes: (i) fine-tuning the teacher model $\mathcal{T}$ on a public auxiliary dataset $\mathcal{D}_{aux}$, and (ii) distilling the student model $\mathcal{S}$ on the same dataset $\mathcal{D}_{aux}$ using the trained teacher model $\mathcal{T}$. Figure 6 illustrates the performance metrics -- balanced accuracy and loss -- of the fine-tuning process on the Tiny-ImageNet dataset. After completing this training, we proceed with distillation for the approximation-enabled student model. Figure 7 shows the performance plot for the distillation process on the Tiny-ImageNet dataset. In the loss behavior plot (on the right), a dashed line marks the start of the second stage. During the first stage, as described in Eq. 16, transformer-layer distillation is performed (showing higher values due to the high-dimensional attention and hidden representation matrices). Then, the second-stage distillation follows, which is a prediction-layer distillation [16]. The learning rate scheduler is employed at epochs $15$ and $25$ with a decay factor of $0.1$. Figures 8 and 9 present analogous performance plots for the Fed-ISIC2019 dataset, using center=0 as the public auxiliary dataset $\mathcal{D}_{aux}$. The learning rate scheduler is employed at epochs $20$ and $40$ with a decay factor of $0.1$.
Figure 6: Performance plot of fine-tuning the teacher model on the public auxiliary dataset ${\mathcal{D}}_{aux}$ (Tiny-Imagenet). The plot on the left shows the balanced accuracy, while the plot on the right presents the loss performance (in a logarithmic scale).
Figure 7: Performance plot of the distillation process on the Tiny-Imagenet dataset. The plot on the left depicts the balanced accuracy across both stages, while the right plot shows the loss performance, with the dashed line indicating the beginning of Stage 2 distillation (Eq. 16).
@@ B.2 Distillation Results
B.3 Approximation Results
Degree $d$ Memory (GB) Latency (s) Attention loss $(\mathcal{L}_{a})$ Representation loss $(\mathcal{L}_{h})$ One-epoch accuracy

### v2

@@ Appendix A Proof of Proposition 1
Given the non-uniqueness property above, an attacker aiming to recover $\bm{A}$}, $\bm{B}$}, and $\bm{C}$} must resort to brute force, testing all possible configurations of $({\bm{A}},{\bm{B}},{\bm{C}})$ that satisfy the products $({\bm{P}},{\bm{Q}},{\bm{R}})$. The total number of combinations of $({\bm{A}},{\bm{B}},{\bm{C}})$ the attacker would need to test is $n!$ and for $n=16$, we need $16!\approx 2\cdot 10^{13}$ combinations.
Appendix B Feature Similarity-based Attack
A potential attack strategy involves matching the $(\mathbf{b}_{\ell},\mathbf{b}_{\ell+1})$ pairs, with $\ell\in[L-1]$, based on their similarity (e.g., L2 distance). After clients decrypt the received permuted representations $\mathbf{b}_{\ell}$, they could attempt to identify approximate nearest representations between the neighboring block outputs, thereby aiding in recovering individual permutation matrices $\Pi_{\ell}$.
Figure 6: Heatmap visualization of the similarity (L2 distance) between corresponding block outputs or hidden representations: (i) between block outputs $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$, and (ii) between block outputs $\mathbf{b}_{11}$ and $\mathbf{b}_{12}$. The X- and Y-axes represent the samples, with a batch size of $8$. An ideal outcome for the attacker would be a perfect minimal diagonal, indicating strong alignment between the representations. While this is not fully observed in the plot, such outcome would reduce the number of brute-force trials required, as outlined in Proposition 1.
To demonstrate this, we provide Figure 6, which illustrates that neighboring blocks exhibit similarity, making it feasible to infer permutation mappings. However, if clients receive randomly sampled blocks instead of sequential ones, the same attack strategy fails (see Figure 7). To mitigate this risk, we introduce stochastic sampling of hidden representations/attention blocks in every forward pass. This ensures that the permutation matrices are never fully exposed, preventing adversaries from reconstructing them (see Proposition 1).
Figure 7: Heatmap visualization of the similarity (L2 distance) between non-consecutive block outputs or hidden representations: (i) between block outputs $\mathbf{b}_{1}$ and $\mathbf{b}_{3}$, and (ii) between block outputs $\mathbf{b}_{9}$ and $\mathbf{b}_{12}$. The X- and Y-axes represent the samples, with a batch size of $8$.
Figure 8: Visualization of the sampling probabilities for $L=12$ transformer blocks under three different scenarios: (i) uniform random sampling with probability $0.5$, (ii) our proposed stochastic sampling strategy, and (iii) stochastic sampling constrained to include either the first or last block.
Additionally, we illustrate the sampling probabilities of transformer blocks under different strategies in Figure 8. Specifically, we compare uniform sampling (included for illustrative purposes), our proposed stochastic sampling strategy depicted in Figure 5, and a constrained version of this stochastic strategy where either the first or last transformer block must always be sampled. The latter strategy ensures meaningful local updates by guaranteeing non-zero input entries for clients. Despite this constraint, the overall expected number of sampled blocks remains unchanged and aligns with the formulation in Eq. 14. Also, the learnable scaling parameter in each adapter layer (see Figure 3), denoted by $\alpha$, addresses potential scaling issues that may arise due to zeroed-out block outputs/hidden representations during stochastic block sampling.
B.1 HE-Compatible Strategy
To maintain privacy while adhering to the computational constraints of homomorphic encryption (HE), we employ the following strategy:
  • Non-selected blocks are set to zeros, ensuring they are not used during the local learning.
  • Due to limited multiplicative depth in HE, we still transmit the non-selected blocks but introduce random noise before sending them to clients.
  • Clients decrypt, re-encrypt, and return the updates. The server then subtracts the added noise before proceeding to the next block, ensuring that adversaries cannot infer the original representations.
This defense mechanism effectively counters feature similarity-based attacks while maintaining the efficiency of our privacy-preserving adaptation framework. A detailed explanation of how sampling is done is provided in Section 4.3.2.
B.2 Complexity Analysis of the Sampling Strategy
Consider a sequence consisting of $L$ consecutive blocks, each assigned either a value of $1$ (sampled) or 0 (not sampled), according to the probabilistic rules depicted in Figure 5:
  • The first block is sampled (assigned the value $1$) with probability $0.5$ and not sampled (assigned 0) with probability $0.5$.
  • For each subsequent block $\ell\geq 2$:
    • If block $\ell-1$ was sampled, then block $\ell$ is deterministically set to 0.
    • If block $\ell-1$ was not sampled, then block $\ell$ is sampled/not sampled with probability $0.5$.
Let $p_{\ell}$ denote the probability of block $\ell$ being sampled. Formally, this probability can be expressed recursively as:
$p_{\ell}=0.5\times(1-p_{\ell-1}),\qquad\text{ with }p_{1}=0.5.$ (11)
Expected Number of Sampled Blocks.
Solving this recurrence relation reveals an equilibrium behavior for large $L$. Specifically, as $L$ gets larger, $p_{\ell}$ converges to a stationary probability $p$, satisfying:
p=0.5(1p).p=0.5(1-p). (12)
Solving this yields the stationary probability $p=\frac{1}{3}$. Therefore, each block has approximately a $\frac{1}{3}$ probability of being sampled. For $L$ blocks, the expected number of sampled blocks $S$ can be expressed as:
$\mathbb{E}[S]=\sum_{\ell=1}^{L}p_{\ell}\approx\frac{L}{3}.$ (13)
When the sampling is repeated for $T$ times, the total expected number of sampled blocks is given by:
$\mathbb{E}[S_{T}]\approx T\times\frac{L}{3}.$ (14)
Appendix C How Does the Distillation Work?
@@
We adopt the transformer distillation approach [24], which consists of two key phases: (i) transformer-layer distillation and (ii) prediction-layer distillation. For generality, let the original transformer model be referred to as the teacher model $(\mathcal{T})$ and the approximation-enabled transformer model as the student model $(\mathcal{S})$. We assume that both models share the same architecture but differ only in their non-linear components (Softmax, GELU, LayerNorm). Further, we outline the primary sub-layers of the transformer blocks:
@@
Figure 9: Schematic illustration of the layer-wise knowledge distillation (Equations 17 and 18).
Datasets Methods Is double-blind? No. of params (M) Latency (ms) Memory (GB)
CIFAR-10 / SVHN Full fine-tuning $82.1096$ (11247.89×)(11247.89\ $\times$) $47.4$ (3.14×)(3.14\ $\times$) $18.13$
LoRA $0.2886$ $(39.54\times)$ $38.5$ $(2.55\times)$ $15.73$
Adapter tuning $3.3933$ $(464.68\times)$ $37.9$ $(2.51\times)$ $14.72$
Linear probing $0.0073$ (1.00×)(1.00\ $\times$) $15.1$ (1.00×)(1.00\ $\times$) $9.39$
BlindFed $0.2536$ (34.74×)(34.74\ $\times$) $15.7$ (1.04×)(1.04\ $\times$) $9.08$
CIFAR-100 Full fine-tuning $82.1756$ (1121.09×)(1121.09\ $\times$) $47.4$ (3.03×)(3.03\ $\times$) $18.13$
LoRA $0.3546$ $(4.84\times)$ $38.4$ $(2.45\times)$ $16.42$
Adapter tuning $3.4593$ $(47.19\times)$ $37.8$ $(2.41\times)$ $15.41$
Linear probing $0.0733$ (1.00×)(1.00\ $\times$) $15.7$ (1.00×)(1.00\ $\times$) $9.40$
BlindFed $0.3196$ (4.36×)(4.36\ $\times$) $16.3$ (1.04×)(1.04\ $\times$) $9.08$
Fed-ISIC2019 Full fine-tuning $82.1082$ (13916.64×)(13916.64\ $\times$) $51.2$ (2.54×)(2.54\ $\times$) $17.32$
LoRA $0.2871$ $(48.66\times)$ $38.5$ $(1.91\times)$ $15.73$
Adapter tuning $3.3919$ $(574.89\times)$ $37.9$ $(1.88\times)$ $14.72$
Linear probing $0.0059$ (1.00×)(1.00\ $\times$) $20.2$ (1.00×)(1.00\ $\times$) $8.71$
BlindFed $0.2522$ (42.75×)(42.75\ $\times$) $21.1$ (1.05×)(1.05\ $\times$) $8.39$
@@
where $h$ is the number of attention heads. And the distillation of the hidden representation is formulated as follows:
@@
Figure 10: Comparison of the efficiency of different adaptation methods in terms of the number of parameters, latency, and memory usage across three datasets: CIFAR-10/SVHN, CIFAR-100, and Fed-ISIC2019. Each plot presents one of the three key metrics, with bars grouped by dataset and color-coded by method. Latency is measured per data point and includes both forward and backward passes. The memory column indicates the GPU memory required for training. Our proposed BlindFed method achieves a favorable balance between efficiency and performance. The results are derived from Table 4.
@@
$\mathcal{L}=\begin{cases}\mathcal{L}_{a}+\mathcal{L}_{h}&\qquad\triangleright\ \text{Stage I},\\ \mathcal{L}_{p}&\qquad\triangleright\ \text{Stage II}.\end{cases}$ (21)
@@
Appendix D Additional Experimental Results
D.1 Baseline Approaches
In this section, we describe the baseline methods used for comparison. Full fine-tuning involves updating all parameters of the pre-trained model by initializing them with pre-trained weights and subsequently training them with gradient updates. In federated experiments, this corresponds to the standard FedAvg algorithm [40]. Linear probing [4] fine-tunes only the classifier head while keeping the backbone frozen. In federated experiments, only the trainable parameters of the head are shared and averaged. LoRA (Low-Rank Adaptation) [21] is a popular parameter-efficient fine-tuning approach, freezing the original weights while training additional low-rank matrices. Its federated counterpart, the FedIT algorithm [65], applies FedAvg specifically to LoRA parameters. Adapter tuning [20] involves inserting lightweight adapter layers between self-attention (attention) and feed-forward modules, connected via residual connections. Each adapter layer comprises two fully connected layers with biases, separated by a nonlinearity (ReLU activation).
Figure 11: Performance plot of fine-tuning the teacher model on the public auxiliary dataset ${\mathcal{D}}_{aux}$ (Tiny-Imagenet). The left plot shows the balanced accuracy, while the right presents the loss performance (log scale).
Figure 12: Performance plot of the distillation process on the Tiny-Imagenet dataset. The left plot depicts the balanced accuracy across both stages, while the right shows the loss performance, with the dashed line indicating the beginning of Stage 2 distillation (Eq. 21).
@@ D.1 Baseline Approaches
D.2 Main Results
Table 4 provides a comprehensive comparison of our method against several baseline approaches, including Full fine-tuning, LoRA, Adapter tuning, and Linear probing, across multiple datasets. This table extends the insights provided in Tables 1 and 2, focusing on three key aspects: the number of trainable parameters, latency, and GPU memory requirements. These results offer a deeper understanding of the trade-offs involved in adapting foundation models for federated learning while maintaining computational feasibility. As observed in the table, full fine-tuning consistently requires a significantly larger number of trainable parameters compared to all other methods, leading to substantially higher memory consumption and latency. In contrast, our method remains efficient, exhibiting a parameter count close to that of LoRA while requiring even less GPU memory.
Notably, across all datasets, our approach remains competitive with linear probing in terms of latency while requiring only a marginal increase in parameter storage. Specifically, in the CIFAR-10/SVHN setting, our method achieves near-optimal latency, operating at only a $4\%$ increase over linear probing, despite introducing a significantly greater degree of adaptability. The reduction in trainable parameters compared to LoRA and Adapter tuning further highlights the effectiveness of our approach in maintaining model compactness without sacrificing computational efficiency.
From a resource perspective, the memory footprint of our method remains consistently low (due to the elimination of the need for backpropagation through the backbone), demonstrating an advantage over more conventional fine-tuning techniques. While full fine-tuning incurs nearly twice the memory cost, our method achieves comparable memory usage to linear probing while outperforming LoRA and Adapter tuning in terms of both efficiency and adaptability.
These findings underscore the scalability and practicality of our method in federated learning applications. By achieving an optimal balance between parameter efficiency, computational cost, and adaptability, our approach effectively mitigates the overhead typically associated with fine-tuning. Moreover, it demonstrates compatibility with homomorphic encryption techniques, achieving lower computational costs compared to conventional fine-tuning methods.
Degree $d$ Memory (GB) Latency (s) Attention loss $(\mathcal{L}_{a})$ Representation loss $(\mathcal{L}_{h})$ One-epoch accuracy $(\%)$